A uniform magnetic field B stores energy of 7.50×10–3J in a box that is 1.66 m3 in volume. What is the strength of the magnetic field B?
Answer #1

As

Energy stored,E=7.5010-3 J

Volume,V=1.66 m3

Energy density is given by

Energy stored/Volume=B2/20

So

B=(2E0/V)1/2

=(27.5010-3410-7/1.66)1/2=1.06610-4 T